a right triangle has hypotenuse c=8 inches and leg b=5 inches find the exact length of the missing side
Answer by TimothyLamb(4379)

h = sqrt( aa + bb )
8 = sqrt( aa + 5*5 )
8 = sqrt( aa + 25 )
8^2 = (sqrt( aa + 25 ))^2
64 = aa + 25
aa = 64 - 25
aa = 39
a = sqrt( 39 )
a = 6.2449979984
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php